# NoTeeline: Supporting Real-Time, Personalized Notetaking with LLM-Enhanced Micronotes

### v2

@@ Abstract.
Taking notes quickly while effectively capturing key information can be challenging, especially when watching videos that present simultaneous visual and auditory streams. Manually taken notes often miss crucial details due to the fast-paced nature of the content, while automatically generated notes fail to incorporate user preferences and discourage active engagement with the content. To address this, we propose an interactive system, NoTeeline, for supporting real-time, personalized notetaking. Given ‘micronotes’, NoTeeline automatically expands them into full-fledged notes using Large Language Model (LLM). The generated notes build on the content of micronotes by adding relevant details while maintaining consistency with the user’s writing style. In a within-subjects study (n=12), we found that NoTeeline creates high-quality notes that capture the essence of their micronotes with 93.2% factual correctness and accurately align with their writing style (8.33% improvement). Using NoTeeline, participants could capture their desired notes with significantly reduced mental effort, writing 47.0% less text and completing their note in 43.9% less time compared to a manual notetaking baseline. Our results suggest that NoTeeline enables users to integrate LLM assistance in a familiar notetaking workflow while ensuring consistency with their preference.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing User interface toolkitsccs: Human-centered computing Natural language interfaces
[…]
MyFrameroundcorner=3pt, linecolor=gray, outerlinewidth=0.5pt, innertopmargin=5pt, innerbottommargin=5pt, innerrightmargin=5pt, innerleftmargin=5pt, backgroundcolor=gray!8!white, frametitleaboveskip=
Figure 1. NoTeeline enables users to take notes while watching videos on-the-fly by automatically converting brief micronotes into fully-expanded notes in their writing style. While watching a video, the user jots down quick keypoints (a). NoTeeline retrieves the transcript at the timestamp (b). In addition, NoTeeline references a small collection of the user’s previous notes (c). Given (b) and (c), NoTeeline automatically expands the micronote into a complete note in their writing style (1). Inspired by the Cornell notetaking system (Saran et al., 2022), NoTeeline generates review questions (2), a brief summary (3), and groups notes into key themes (4) to assist users in note review.
@@ 1. Introduction
Effective notetaking is crucial for capturing and organizing information, allowing users to record essential details for later review. Whether in academic, professional, or personal settings, notes help individuals track important points, revisit information when needed, and maintain focus on the task at hand without getting overwhelmed by the flow of information. However, notetaking can be very challenging in a fast-paced environment, such as when watching information-dense videos or videos that are difficult to pause. Users may struggle with the high mental demands, forcing them to divide their attention between absorbing new content and documenting it (Piolat et al., 2005; Jansen et al., 2017; Piolat, 2007). For example, when taking notes from information-dense videos, people often pause the video frequently in order to write notes, which disrupts the natural flow of content engagement (Cao et al., 2022) and forces the user to dedicate mental and physical effort to the mechanical act of notetaking. The inability to pause (e.g. in some live-streamed content or content on platforms that lack full media controls) adds even more pressure, often resulting in incomplete notes that may be difficult to interpret later (McCoy, 2020; Kafipour and Jafari, 2021; Flanigan and Titsworth, 2020)
@@
With the emergence of Large Language Models (LLMs,) people have explored building automated notetaking tools, such as NoteGPT (NoteGPT.io, 2024; Notion, 2024; Otter.ai, 2023). These tools typically work by automatically generating notes or summaries from the video transcript or live lecture recordings in one go. Thus, the generated notes lack personalization (Biermann et al., 2022). For example, the notes may be too general and not tailored to the user’s specific needs. A user may only be interested in specific topics of the material, or they might prefer their notes to be written in a particular style that reflects how they process information. Automated notes can feel impersonal and robotic (Biermann et al., 2022).
Prior studies have observed that, in fast-paced notetaking situations, people often resort to shorthand techniques like micronotes(Lin et al., 2004). For example, they may jot down short keypoints containing shortened words or informational notations in order to write things down more quickly (see Figure 1a). However, while micronotes help users keep up with the pace, they are often too brief or vague to be useful during the later review. As a result, these micronotes are often abandoned and not revisited (Kleek et al., 2009). Thus, we observed a gap between quickly capturing information and having detailed, meaningful notes that can be revisited with clarity.
In this work, we leverage the practice of micro-notetaking and build an assistive tool around this practice, specifically for videos. Video content presents a unique challenge for notetaking due to its temporal nature and often being information-dense, which leaves little time for users to capture important details. We introduce NoTeeline  [^1] NoTeeline’s name is inspired by Teeline(Contributors, 2024), a popular shorthand technique., an interactive notetaking assistance tool that aims to automatically expand user micronotes into complete notes in real-time that also aligns with their personal writing style. We propose a novel pipeline that leverages a window of the video transcript as well as a handful of examples of the user’s previous notes. Through a within-subjects user study, we found that NoTeeline helps users effectively capture notes with minimal disruption (e.g., reduced video pauses), capturing more content using fewer words and shorter time compared to a manual notetaking baseline. Participants rate the generated note expansions as high quality (average 6.0 on 7-point Likert scale, higher is better), align with the participants’ micronote content (relatively high semantic similarity between micronotes and expanded notes measured with SBERT score (Reimers and Gurevych, 2019)), are consistent with participants’ writing style (8.33% improvement in chi-squared Distance (Kilgarriff, 2001)), and the notes contain low hallucination (93.2% factual correctness score measured with HHEM score (Vectara, 2024)).
In summary, this work makes the following contributions.
@@
    We designed NoTeeline, an interactive notetaking tool built around the concept of micronotes. NoTeeline helps users write quick notes in fast-paced contexts and utilizes LLMs to expand micronotes into full-fledged notes.
@@
    We conducted a within-subjects user study to evaluate the performance and utility of NoTeeline. Our quantitative and qualitative results show that with NoTeeline, users can take high-quality notes with higher efficiency and satisfaction, compared to a manual notetaking baseline.
@@ 4.1. Procedure
We conducted a needfinding study to understand the current challenges users encounter with their current notetaking process. Specifically, we want to uncover underlying factors that make writing the full note difficult. The results of this study informed the design choices we made for NoTeeline. Each session was approximately 20-30 minutes long.
Our participants reported using both manual notetaking tools as well as AI-assisted tools. The most common app used by participants was the built-in Macbook Notes app ($n=9$). Many participants use pen-and-paper notes ($n=7$) and later digitize them manually. In terms of AI-based tools, some participants used ChatGPT ($n=4$) for summarizing meetings and video lectures. Others used Notion AI, Figjam AI, and Gemini plugin in Google Docs for specific use cases.
@@ 4.3. Design Goals
D2. Maintain consistency with user’s writing style and preferences to ensure user agency. The system should take into account the specific writing styles of the user and remain consistent with their personal style. At the same time, the final note should exclude information users are not interested in and should not include any hallucinated information (C2).
@@ 5.1. Interface Components
5.1.2. Notetaking Layout
. To better situate the interface around an effective workflow, we took inspiration from the Cornell method to construct the layout (Pauk and Owens, 2010; Saran et al., 2022; Piolat, 2007). We have three main regions – 1) Notes: the region for writing and organizing notes; 2) Cues: auto-generated questions to review the note; and 3) Summary: main takeaway of the note.
@@ 5.2. System Walkthrough
Feeling content with the organization of her notes, Kate clicks on [Uncaptioned image] to generate a personalized TL; DR summary for her conclusion slide. Finally, Kate wants to anticipate potential questions from her audience. She opens the [Uncaptioned image] panel. Attempting to answer from memory, Kate correctly addresses the first question but stumbles on the second, prompting her to revisit her notes for a quick recap. Satisfied with her understanding and the comprehensive coverage of her topic, she confidently moves on to creating her presentation using her notes from NoTeeline.
6. Prototype Implementation
NoTeeline is implemented using React.js and is hosted as a web app. All LLM-based operations and text manipulation are performed using the OpenAI model, GPT-4-Turbo with hyperparameters set to temperature = 0.5, seed = 1. We chose GPT-4 over other LLMs due to its high-quality output and support for a longer size context window, especially suitable for longer videos (OpenAI, 2023). We store all the data locally using Chrome localstorage.
@@ 7. User Evaluation
Participants. We invited 12 participants from our needfinding interviews to participate in a user study. Participants used both NoTeeline and a baseline to capture notes and evaluate their performance. They were divided into four subgroups to counterbalance and mitigate any ordering bias. Table 3 shows the distribution of participants across these subgroups.
@@ 8.2. NoTeeline Generates High Quality Note Expansions (RQ2)
During the user study, participants evaluate the quality of the automatically generated full note. On average, users rated the expanded note, $\mu=6$ ($\sigma=1.044$) on a scale of 1 to 7 (higher is better) (Figure 10, 1st row).
@@
We report the linguistic quality (in terms of grammaticality, non-redundancy, focus) of notes using GRUEN (Zhu and Bhat, 2020). The result (Table 4 in Appendix) indicates that the points written using NoTeeline ($\mu=0.84,\sigma=0.07$) are of higher quality (i.e: less grammatical error, less redundant and more focused) than the baseline ( $\mu=0.56,\sigma=0.17$) ($p<0.0001,t(11)=10.87$).
@@ 8.2.3. Correctness
. To ensure our automatically expanded notes are indeed correctly extracting the information from the transcript and not hallucinating any misinformation (C2), we compute the factual correctness of the generated note using HHEM (Vectara, 2024; Zhang et al., 2023; Azamfirei et al., 2023). We noticed a Hallucination Rate = $6.8\%$ and a Factual Consistency Rate = $93.2\%$. This indicates that our generated notes are correct and consistent with the video content $93.2\%$ of the time.
@@ 8.2.7. Error Analysis
Error in LLM generation. GPT-4 sometimes fails to generate the requested note. For example, P7 wrote a micronote to note down her thoughts that are not directly discussed in the video. The model replied with - ‘Please provide the transcript related to the keypoint so I can assist you in creating a note.’. The occurrence of such error is 0.72% (4 out of 559 total calls).
@@ 11. Conclusion
Active notetaking from videos remains challenging. While existing tools can automatically generate notes, they do not necessarily capture the details and stylistic nuances users care about. Our needfinding study reveals that users often do not prefer automatically generated notes due to concerns about accuracy and lack of customization. On the other hand, NoTeeline leverages rich contextual information to generate more accurate and more personalized notes. Our user study reveals that NoTeeline helps write notes faster with $93.2\%$ correctness and achieves closer stylistic characteristics to users’ handwritten notes. We demonstrate the utility of NoTeeline in a variety of usage scenarios users reported and provide design implications for future works in AI-assisted notetaking systems.
@@
By helping users take notes faster, NoTeeline has the potential to make notetaking more accessible for individuals who might otherwise struggle with factors such as language fluency (e.g., as demonstrated by P7’s comments about difficulties of being a non-native English speaker) and physical disabilities. While there exist arguments about whether over-reliance on AI-generated content may lead to passive engagement or diminish personal organizational skills, we show how our tool exists to amplify, not override, user engagement in the generation process. We also expect this to decrease the risk of AI biases that might otherwise occur in fully autogenerated notes without any sort of user engagement.
Acknowledgements

### v3

@@ Abstract.
Taking notes quickly while effectively capturing key information can be challenging, especially when watching videos that present simultaneous visual and auditory streams. Manually taken notes often miss crucial details due to the fast-paced nature of the content, while automatically generated notes fail to incorporate user preferences and discourage active engagement with the content. To address this, we propose an interactive system, NoTeeline, for supporting real-time, personalized notetaking. Given ‘micronotes’, NoTeeline automatically expands them into full-fledged notes using a Large Language Model (LLM). The generated notes build on the content of micronotes by adding relevant details while maintaining consistency with the user’s writing style. In a within-subjects study (n=12), we found that NoTeeline creates high-quality notes that capture the essence of participant micronotes with 93.2% factual correctness and accurately align with participant writing style (8.33% improvement). Using NoTeeline, participants could capture their desired notes with significantly reduced mental effort, writing 47.0% less text and completing their notes in 43.9% less time compared to a manual notetaking baseline. Our results suggest that NoTeeline enables users to integrate LLM assistance in a familiar notetaking workflow while ensuring consistency with their preferences—providing an example of how to address broader challenges in designing AI-assisted tools to augment human capabilities without compromising user autonomy and personalization.
@@
copyright: acmlicensedjournalyear: 2025copyright: ccconference: 30th International Conference on Intelligent User Interfaces; March 24–27, 2025; Cagliari, Italybooktitle: 30th International Conference on Intelligent User Interfaces (IUI ’25), March 24–27, 2025, Cagliari, Italydoi: 10.1145/3708359.3712086isbn: 979-8-4007-1306-4/25/03ccs: Human-centered computing User interface toolkitsccs: Human-centered computing Natural language interfaces
Figure 1. NoTeeline facilitates on-the-fly notetaking while watching videos by expanding user-defined micronotes into richer-context notes written in their writing style. While watching a video, the user jots down quick keypoints (a). NoTeeline retrieves the transcript at the timestamp (b). In addition, NoTeeline references a small collection of the user’s previous notes (c). Given (b) and (c), NoTeeline automatically expands the micronote into a complete note in their writing style (1). Inspired by the Cornell notetaking system (Saran et al., 2022), NoTeeline generates review questions (2), a brief summary (3), and groups notes into key themes (4) to assist users in note review.
@@ 1. Introduction
Effective notetaking is crucial for capturing and organizing information. Whether in academic, professional, or personal settings, notes help individuals track important points, revisit information when needed, and maintain focus on the task at hand without getting overwhelmed by the flow of information. However, notetaking can be particularly challenging in a fast-paced environment like watching information-dense videos. Users may struggle with the high mental demands, forcing them to divide their attention between absorbing new content and documenting it (Piolat et al., 2005; Jansen et al., 2017; Piolat, 2007). Users frequently pause the video to write notes, disrupting the natural flow of content engagement (Cao et al., 2022) while forcing them to dedicate mental and physical effort to the mechanical act of notetaking. The inability to pause (e.g. in some live-streamed content or content on platforms that lack full media controls) adds even more pressure, often resulting in incomplete notes that may be difficult to interpret later (McCoy, 2020; Kafipour and Jafari, 2021; Flanigan and Titsworth, 2020).
@@
With the emergence of Large Language Models (LLMs,) there has been significant interest in developing automated notetaking tools from video transcripts (e.g. NoteGPT (NoteGPT.io, 2024), Otter (Otter.ai, 2023)). While these tools may simplify notetaking, one common limitation is their lack of personalization. Automatically generated notes are often generic and fail to address individual user needs: for instance, users may have specific writing style preferences (Biermann et al., 2022) or specific topics of interest. As a result, these notes can feel impersonal and robotic, diminishing their usefulness for everyday use (Biermann et al., 2022).
Alternatively, users may adopt shorthand notetaking strategies (i.e. micronotes) to manage the high mental demand of fast-paced information consumption (Lin et al., 2004). They may jot down short keypoints containing shortened words or informational notations in order to write things down more quickly (see Figure 1a). While micronotes help users keep up with the pace, they are often too vague to be useful during the later review. As a result, micronotes are often abandoned afterwards and not revisited (Kleek et al., 2009). Thus, we observed a gap between quickly jotting down notes to capture information and having meaningful notes that can be revisited with clarity.
In this work, we leverage the practice of micro-notetaking and build an assistive note taking tool for videos. We introduce NoTeeline  [^1] NoTeeline’s name is inspired by Teeline(Contributors, 2024), a popular shorthand technique., an interactive notetaking assistance tool that aims to automatically expand user micronotes into complete notes in real-time, generated using their personal writing style. We propose a novel pipeline that leverages video transcript for information extraction and a small handful of the user’s previous note examples for preference modeling. Through a within-subjects user study, we found that NoTeeline helps users effectively capture notes with minimal disruption (e.g., reduced video pauses), capturing more content using fewer words and shorter time compared to a manual notetaking baseline. Participants rate the generated note expansions as high quality (average 6.0 on 7-point Likert scale, higher is better), align with the participants’ micronote content (relatively high semantic similarity between micronotes and expanded notes measured with SBERT score (Reimers and Gurevych, 2019)), are consistent with participants’ writing style (8.33% improvement in chi-squared distance (Kilgarriff, 2001)), and the notes contain low hallucination (93.2% factual correctness score measured with HHEM score (Vectara, 2024)). For reproducibility, we provide our source code and user logs at: https://github.com/oaishi/NoTeeline
In summary, this work makes the following contributions:
@@
    We introduce NoTeeline, a proof-of-concept notetaking tool built around the existing practice of micronotaking. NoTeeline leverages LLM pipeline to expand rough micronotes into full-fledged notes. NoTeeline facilitates a fundamentally different approach to notetaking, centered on taking quick micronotes and letting LLM fill in the details, saving mental and physical effort.
@@
    We conducted a within-subjects user study to evaluate the performance and utility of NoTeeline. Our quantitative and qualitative results show that with NoTeeline, users can take high-quality notes with higher efficiency, reduced disruption, and higher overall satisfaction, compared to a manual notetaking baseline. Users rated notes created using NoTeeline with high consistency to their writing style (8.33% improvement in chi-squared distance via SBERT score) and low hallucination (93.2% HHEM score).
@@ 4.1. Procedure
Few studies have explored how people use LLM-based note-taking systems alongside manually written notes. To better understand this area, including overall user perception and limitations, we conducted a needfinding study. Please refer to Appendix F for the complete list of study questions. Each study session was approximately 20-30 minutes long. Table 1 shows the details for the participants who took part in the study. Our participants are in the age range of 20-30 ($\mu=24.83,\sigma=2.37$)
The most common app used by participants was the built-in Macbook Notes app ($n=9$). Many participants use pen-and-paper notes ($n=7$) and later digitize them manually. In terms of AI-based tools, some participants used ChatGPT ($n=4$) for summarizing meetings and video lectures. Others used Notion AI, Figjam AI, and Gemini plugin in Google Docs for specific use cases.
ID Nationality Age Gender Occupation
P1 South Korean 30 Male Graduate Student (CS)
P2 Chinese 25 Female Undergraduate (HCI)
P3 Chinese 25 Female UX Designer
P4 American 22 Female
Undergraduate
(Information Systems + HCI)
P5 Indian 25 Female Graduate Student (HCI)
P6 American 21 Female
Undergraduate
(Information Systems)
P7 South Korean 25 Female
Graduate student
(CS + Design)
P8 Indian 25 Female
Graduate Student
(Robotics + CS)
P9 Singaporean 24 Male
Undergraduate
(Fine Arts + HCI)
P10 American 22 Female
Undergraduate
(Civil Engr.)
P11 Indian 27 Male Product Manager
P12 American 27 Male
Software Engineer
Table 1. User study participant demographics.
@@ 4.3. Design Goals
D2. Allowing users to guide note generation using micronotes. The system should enable users to guide note generation using micronotes, directing the LLM to focus on relevant video segments. LLM-expanded notes should exclude irrelevant or erroneous information (C2). Additionally, the system should adapt to each user’s writing style, maintaining consistency.
@@ 5. NoTeeline
[…]
The NoTeeline user interface consists of (a) the Cues panel (outlined in blue) for review questions generated based on the notes, (b) the Notes panel (outlined in orange) for user notetaking and note expansion, and (c) the Summary panel (outlined in purple) for note summary. The layout is inspired by the Cornell method, as seen in the side-by-side comparison. The onboarding session, highlighted on the left, collects example notes from the user.
@@ 5.1. Interface Components
5.1.2. Notetaking Layout.
To better situate the interface around an effective workflow, we took inspiration from the Cornell method to construct the layout (Saran et al., 2022). We have three main regions – 1) Notes: the region for writing and organizing notes; 2) Cues: auto-generated questions to review the note; and 3) Summary: main takeaway of the note.
@@
[…]
Micronote expansion. Users can toggle between the micronote and the expanded note by clicking the Expand/Reduce buttons.
@@
[…]
Note organization by themes (some texts have been cropped for clarity). Users can organize the notes by themes or toggle back to the original serial in which they took the notes.
@@ 5.2. System Walkthrough
Feeling content with the organization of her notes, Kate clicks on [Uncaptioned image] to generate a personalized TL; DR summary for her conclusion. Finally, Kate wants to anticipate potential questions from her audience. She opens the [Uncaptioned image] panel. Attempting to answer from memory, Kate correctly addresses the first question but stumbles on the second, prompting her to revisit her notes for a quick recap. Satisfied with her understanding and the comprehensive coverage of her topic, she confidently moves on to creating her presentation using her notes from NoTeeline.
5.3. Prototype Implementation
NoTeeline is implemented as a web app using React.js and Chakra UI. A NodeJS server handles the API requests to the LLM. We used youtube-transcript-api, a Python package to extract the transcript from youtube videos. LLM-based operations and text manipulation are performed using the OpenAI model, GPT-4-Turbo with hyperparameters set to temperature = 0.5, seed = 1. To reduce the randomness in API response, we used fixed model checkpoints with a predefined seed value. We chose GPT-4 over other LLMs due to its high-quality output and support for a longer context window, especially suitable for longer videos (OpenAI, 2023). On average, OpenAI takes $2.00$ seconds to expand all the micronotes in a single video (Please refer to table 8 to view more details). We store all the data (e.g. note points, transcription, summary ar onboarding examples) in Chrome localstorage.
@@ 6. User Evaluation
Participants. We invited 12 participants from our needfinding interviews to participate in a user study. Participants used both NoTeeline and a baseline to capture notes and evaluate their performance. They were divided into four subgroups to counterbalance and mitigate any ordering bias. Table 4 shows the distribution of participants across these subgroups. Participants were not exposed to information regarding the NoTeeline system before the user study.
@@ 7. Results and Findings
[…]
Distribution of pause and skips used by each participant in NoTeeline (left) and Baseline (right). In NoTeeline, the count of pause and skip is significantly less than the baseline. An empty bar depicts that the user did not pause or skip at all.
@@ 7.1. NoTeeline Helps Users Write Notes Efficiently with Reduced Disruption (RQ1)
[…]
Comparison of (a) note length, (b) writing time and (c) count for each participant between NoTeeline and baseline. The majority of the users took the higher amount of notes in NoTeeline in less time.
@@
[…]
User study results (N=12). Boxplots from left to right: (a) workload measured with NASA TLX (Hart and Staveland, 1988) (7-point Likert scale, lower is better), (b) usability measured with SUS (Brooke et al., 1996) (7-point Likert scale, higher is better)
@@ 7.2. NoTeeline Generates High Quality Note Expansions (RQ2)
During the user study, participants evaluated the quality of the automatically generated full note. Eleven out of 12 participants preferred the NoTeeline expanded note and summary compared to the automatically generated ones (Figure 10).
@@
To measure linguistic quality and grammatical correctness, we report the GRUEN (Zhu and Bhat, 2020) of the expanded notes. The result (Table 5 in Appendix) indicates that the points written using NoTeeline ($\mu=0.84,\sigma=0.07$) are of higher quality (i.e: less grammatical error, less redundant and more focused) than the baseline ( $\mu=0.56,\sigma=0.17$) ($p<0.0001,t(11)=10.87$).
@@ 7.2.3. Correctness
To ensure our automatically expanded notes are indeed correctly extracting the information from the transcript and not hallucinating any misinformation (C2), we compute the factual correctness of the generated note using HHEM (Vectara, 2024; Zhang et al., 2023; Azamfirei et al., 2023). We noticed a Hallucination Rate = $6.8\%$ and a Factual Consistency Rate = $93.2\%$. This indicates that our generated notes are correct and consistent with the video content $93.2\%$ of the time.
@@ 7.2.4. Consistency with Personal Writing style
[…]
Chi-squared distance(Kilgarriff, 2001) between the expanded notes from micronote and user’s manual note. Examples of previous note collected during onboarding help us to achieve a better style similarity with the user’s writing. Green and Red denote positive and negative relative performance change, respectively.
@@ 7.2.7. Error Analysis
GPT-4 sometimes fails to generate the requested note. For example, P7 wrote a micronote to note down her thoughts that are not directly discussed in the video. The model replied with - ‘Please provide the transcript related to the keypoint so I can assist you in creating a note.’. The occurrence of such error is 0.72% (4 out of 559 total calls).
@@ 7.3.3. Comments on NoTeeline’s Individual Features
[…]
Participants’ responses to utility and usability of NoTeeline interface and various features (7-point Likert scale, higher is better).
@@ 10. Conclusion
Active notetaking from videos remains challenging. While existing tools can automatically generate notes, they do not necessarily capture the details and stylistic nuances users care about. Our needfinding study reveals that users often do not prefer automatically generated notes due to concerns about accuracy and lack of customization. NoTeeline leverages rich contextual information to generate more accurate and more personalized notes. Our user study reveals that NoTeeline helps write notes faster with $93.2\%$ correctness and achieves closer stylistic characteristics to users’ handwritten notes. We demonstrate the utility of NoTeeline in a variety of usage scenarios users reported and provide design implications for future works in AI-assisted notetaking systems.
@@
By helping users take notes faster, NoTeeline has the potential to make notetaking more accessible for individuals who might otherwise struggle with factors such as language fluency (e.g., P7’s comments about difficulties of being a non-native English speaker) and physical disabilities. While there exist arguments about whether over-reliance on AI-generated content may lead to passive engagement or diminish personal organizational skills, we show how our tool exists to amplify, not override, user engagement in the generation process. We also expect this to decrease the risk of AI biases that might otherwise occur in fully autogenerated notes without any sort of user engagement.
Acknowledgements.